# Efficient Neuro-Symbolic Learning of Constraints & Objective

### v1

@@ Abstract.
Background: In the ongoing quest for hybridizing discrete reasoning with neural nets, there is an increasing interest in neural architectures that can learn how to solve discrete reasoning or optimization problems from natural inputs, a task that Large Language Models seem to struggle with.
@@
Methods: Our new probabilistic loss allows for learning both the constraints and the objective, thus delivering a complete model that can be scrutinized and completed with side constraints. By pushing the combinatorial solver out of the training loop, our architecture also offers scalable training while exact inference gives access to maximum accuracy.
Results: We empirically show that it can efficiently learn how to solve NP-hard reasoning problems from natural inputs. On three variants of the Sudoku benchmark – symbolic, visual, and many-solution –, our approach requires a fraction of training time of other hybrid methods. On a visual Min-Cut/Max-cut task, it optimizes the regret better than a Decision-Focused-Learning regret-dedicated loss. Finally, it efficiently learns the energy optimization formulation of the large real-world problem of designing proteins.
@@ 1. Introduction
In recent years, several hybrid neural architectures have been proposed that integrate discrete reasoning or optimization within neural network architectures. The motivations for these works are varied. For discrete optimization and reasoning experts, the main motivation is to deal with uncertainty proactively. Available data is leveraged to produce a solution (decision), whose quality approaches that achievable with a full knowledge of the problem to solve. This motivation has led to the introduction of Decision-Focused Learning (DFL) (Mandi et al., 2024). Such hybrid architectures are also of interest for users who have a dataset of discrete structured objects of interest (sequences, trees, graphs,…) with associated correlated features and want to be able to predict or generate similar structured objects given only their features. Finally, architectures combining discrete optimisation and learning fit within the neuro-symbolic framework. It gathers approaches capable of processing natural inputs – such as images, text or molecules – while simultaneously exhibiting extensive logical reasoning capabilities. Given the accumulating evidence that large language models are limited in their ability to perform logical reasoning (Bachmann and Nagarajan, 2024; Ye et al., 2025; Lin et al., 2025; Zubic et al., 2025), neuro-symbolic systems are a promising alternative research direction in the quest for sparks-of-AGI (artificial general intelligence) (Ye et al., 2025).
@@
In machine learning, the problem of predicting complex discrete objects (sequences, trees, graphs, …) has been addressed in the field of Structured Output Prediction (SOP): given a training set of pairs $(\omega^{\ell},\mathbf{y}^{\ell})$, where $\omega^{\ell}$ is a suitable input data and $\mathbf{y}^{\ell}$ is a corresponding structured output, the aim is to predict a new high-quality structured output $\mathbf{y}$ from new input data $\omega$. In this setting, $\mathbf{y}$ can be predicted by arbitrary methods. In this paper, we are more specifically interested in architectures where $\mathbf{y}$ is an optimal solution of a predicted combinatorial optimization (CO) problem instance with unknown objective function $F$ with parameters $\mathbf{c}$ and possibly unknown constraints $\mathcal{C}$. In a generative context, without retraining, the predicted CO instance can be arbitrarily constrained to produce optimal $\mathbf{y}$ that satisfy additional properties. Alternatively, additional objectives can be considered to produce optimal compromises.
@@
This setting raises several challenges. First, the discrete nature of variables defines zero gradients almost everywhere, preventing direct learning by gradient-based optimisation. Second, predicting optimal solutions of CO problems using exact solvers or heuristic algorithms may be computationally demanding and is not autograd-differentiable. The computational complexity is specifically problematic at training time, with massive training sets that may contain large NP-hard problem instances. We designed our approach to scale to such demanding situations for broad applicability. Our approach can also deal with partially observable settings, where observed solutions $\mathbf{y}^{\ell}$ may contain the values of only a subset of all variables.
@@
Our contributions are threefold: we provide an extensive review of the literature, we introduce a neuro-symbolic architecture combining arbitrary deep learning layers with a final discrete NP-hard Graphical Model (GM) reasoning layer, and propose a new loss function that efficiently deals with logical information presented as constraints. We use discrete GMs (Cooper et al., 2020) as the reasoning language because GMs have been used to represent both Boolean functions or constraints (in propositional logic, constraint networks) and numerical functions or objectives (in partial weighted Max-SAT, cost function networks). For learning, a probabilistic interpretation of such models is available in stochastic GMs such as Markov Random Fields (MRFs), where infinite costs represent zero probabilities (infeasibility).
@@
We exploit this probabilistic interpretation and target the optimization of the negative loglikelihood (NLL) of the training set. Because of its intractability, we use an asymptotically consistent and scalable approximation of the NLL known as the Negative Pseudo-LogLikelihood (NPLL) (Besag, 1975). We observe that the NPLL may become incapable of estimating many large costs (Montanari and Pereira, 2009) and, therefore, of learning constraints. We analyze the reasons for this limitation and propose the E-PLL, which combines NPLL scalability with the ability to learn constraints. At inference, the neural architecture receives natural inputs and outputs a fully-parameterized discrete GM. This model can then be solved using any GM optimization solver to produce a solution.
@@
With this differentiable informative loss, our architecture efficiently learns constraints from natural inputs to, e.g., solve the Visual Sudoku (Wang et al., 2019; Brouard et al., 2020). Due to its ability to deal with unobserved variables, it can solve the associated symbol grounding problem without cheating (Chang et al., 2020). The use of an exact prover during inference provides 100% accuracy on hard symbolic Sudoku instances, even with small training sets. The same accuracy is obtained on a many-solution Sudoku data set (Nandwani et al., 2021) and more complex Futoshiki grids (Şen and Diner, 2024). To show that our approach applies in a DFL context where constraints are known, we also tackle Min-Cut and Max-Cut instances where edge weights are represented as images. Finally, as a certificate of scalability, we train our architecture on a set of Protein Design instances with more than 1,000 variables.
@@ 2. Related works
Inverse optimization and constraint acquisition.
The idea of learning the parameters of optimization problems from a set of solutions has been explored in various fields. Classical inverse optimization (Chan et al., 2023) starts from a dataset of solutions $(\mathbf{y}^{\ell})$, each being an exact optimum $\mathbf{y}^{\ell}\in\operatorname*{arg\,min}F(\mathbf{y},\mathbf{c})$ with unknown parameters $\mathbf{c}$, that must be identified. Similarly, in constraint programming, constraint acquisition starts from sets of solutions, and possibly non-solutions, of an unknown constraint satisfaction (Beldiceanu and Simonis, 2016; Bessiere et al., 2023) or satisfiability (Kumar et al., 2023) problem that needs to be identified. Compared to our setting, these approaches do not exploit arbitrary observable features $\omega$. In data-driven inverse optimisation (Chan et al., 2023), the optimization problem parameters $\mathbf{c}$ are split into observable and unobservable parameters $\mathbf{c}=\mathbf{c}_{o}\cup\mathbf{c}_{u}$. Starting from a dataset of pairs $(\mathbf{y}^{\ell},\mathbf{c}_{o}^{\ell})$, the aim is to predict $\mathbf{c}_{u}$ so that on average, empirically, each $\mathbf{y}^{\ell}$ is close, in distance or cost, to an optimal solution of $F(\mathbf{y},\mathbf{c}_{o}^{\ell}\cup\mathbf{c}_{o})$. This can be seen as a specific case of our setting, where the observed data $\omega$ is the exact observation of a fraction of the parameters instead of features known to be correlated with the complete problem definition (i.e., both constraints and objective). However, existing approaches are dedicated to continuous (linear, convex, or conic) optimisation problems.
@@
Deep learning has been combined with logic programming into several tools such as DeepProbLog (Manhaeve et al., 2018) or NeurASP (Yang et al., 2020), where deep learning is in charge of perception and Prolog/ASP is in charge of reasoning. But training is restricted to neural network under known constraints on the output.
@@ 2.1. Decision-focused learning
In many discrete optimisation under uncertainty problems, the target output $\mathbf{y}$ is the solution of a CO problem defined by a fixed set of constraints $\mathcal{C}$ and an objective function $F$ with unknown parameters $\mathbf{c}$. The parameters $\mathbf{c}$ can be predicted from historic data of features $\omega$, using a training set of pairs $(\omega^{\ell},\mathbf{c}^{\ell})$. In this setting, Predict-then-Optimize – or Prediction-focused-learning – methods (Mandi et al., 2024) learn to directly predict parameters $\hat{\mathbf{c}}$ from input features $\omega$ with a typical regression loss such as MSE, regardless of the downstream optimisation task. At inference, the fixed discrete model is solved with the predicted $\hat{\mathbf{c}}$. Instead, DFL approaches aim to predict parameters leading to “good” decisions that empirically optimize the average regret, the loss in objective caused by optimizing with $\hat{\mathbf{c}}$ instead of the true parameters $\mathbf{c}$.
DFL differs from our setting by the very nature of the training data. Still, since $\mathcal{C}$ is fixed, DFL training-set pairs $(\omega^{\ell},\mathbf{c}^{\ell})$ can be transformed into pairs $(\omega^{\ell},\mathbf{y}^{\ell})$ where $\mathbf{y}^{\ell}$ satisfies the constraints in $\mathcal{C}$ and optimizes $\mathbf{c}^{\ell}$, just by solving the optimization problem defined by $\mathcal{C}$ and $\mathbf{c}^{\ell}$.
Thus, methods designed for SOP directly apply to DFL settings and have been de facto considered as DFL tools in recent reviews (Mandi et al., 2024). Existing methods have been classified in four families (Mandi et al., 2024): analytical differentiation of optimization mappings, analytical smoothing of optimization mappings, smoothing by random perturbations, and differentiation of surrogate loss functions. Differentiable optimization layers have mostly been developed for convex or conic continuous optimization problems. They can be applied to discrete optimization problems after proper smoothing. This is, for instance, the case of SAT-Net (Wang et al., 2019), which relaxes the discrete MaxSAT problem into a convex semi-definite programming problem. SAT-Net has been used to learn how to solve easy Sudoku and Visual Sudoku (from grid images). It has been pointed out that SAT-Net was ineffective if labels for observed digit images were not provided in the supervised training set (Chang et al., 2020).
@@
Analytical smoothing is also needed for learning linear programming (LPs) problems because of the intrinsically discrete nature of polytope’s vertices. One standard smoothing is the addition of an $L_{2}$ regularization on solutions (Wilder et al., 2019; Niculae and Martins, 2020). For integer LPs (ILPs), it is possible to relax them to LPs (Wilder et al., 2019), possibly with cutting plane strengthening (Ferber et al., 2020) for a better polytope approximation. Discrete (Boolean/One Hot Encoded) pairwise GMs, the formalism we use in this paper, are intrinsically quadratic in their discrete decision variables and linear in their parameters. They can be optimized by reduction to an 01LP, requiring the introduction of a quadratic number of variables. This 01LP model, relaxed to continuous variables, is known as the GM’s “local polytope” (Cooper et al., 2020). Existing approaches for learning ILP could, therefore, be used to learn GMs. Due to their size, the associated (I)LP reductions are challenging to solve (Hurley et al., 2016; Allouche et al., 2014).
@@
Besides smoothing by relaxation or regularization, e.g.  in Fenchel-Young losses (Blondel et al., 2020), smoothing can be achieved by random perturbations (Berthet et al., 2020). These approaches rely on a probabilistic interpretation of the linear objective function using maximum-entropy exponential distributions. For discrete models, assuming minimisation, this usually means Gibbs/Boltzmann distributions with the probability $p(\mathbf{y})\propto\exp(-\langle\mathbf{c},\mathbf{y}\rangle)$. In arbitrary discrete settings, the resulting probability distributions suffer from the intractability of computing the normalizing constant $Z$. This can sometimes be simplified using the Gumbel trick (Huijben et al., 2022), which allows for estimating $Z$ using a series of calls to an optimisation oracle on a Gumbel-perturbed objective parameters $\mathbf{c}$ (Niepert et al., 2021) or using other perturbations (Berthet et al., 2020). These approaches can be interpreted as Fenchel-Young losses (Blondel et al., 2020; Berthet et al., 2020), where the perturbed solutions provide the regularization. The stochastic variants of GMs (Cooper et al., 2020; Koller and Friedman, 2009) we rely on, Markov Random Fields, are decomposable log-linear models that rely on the same family of exponential distributions. Their log-likelihood is intractable for the same reason. It is known to be a Fenchel-Young loss (Blondel et al., 2020), often referred to as the CRF (Conditional Random Field) loss. We instead use and improve a tractable approximation of log-likelihood that exploits the decomposability of GMs.
@@
These approaches try to optimize regret indirectly. The SPO+ loss (Elmachtoub and Grigas, 2022) is an upper bound of regret with a gradient that resembles perturbed optimizers’ gradients. A recent related approach, based on Noise Contrastive Estimation (Mulamba et al., 2021), exploits an incrementally built pool of feasible solutions in a probabilistic interpretation, as above.

### v3

# Scaling Neuro-symbolic Problem Solving: Solver-Free Learning of Constraints and Objectives

@@ Abstract.
Background: In the ongoing quest for hybridizing discrete reasoning with neural nets, there is an increasing interest in neural architectures that can learn how to solve discrete reasoning or optimisation problems from natural inputs, a task that Large Language Models seem to struggle with.
@@
Methods: Our new probabilistic loss allows for learning both the constraints and the objective – possibly non-linear – of a combinatorial problem. Thus, it delivers a complete model that can be scrutinized and completed with side constraints. By pushing the combinatorial solver out of the training loop, our architecture also offers scalable training while exact inference gives access to maximum accuracy.
Results: We empirically show that it can efficiently learn how to solve NP-hard reasoning problems from natural inputs. On three variants of the Sudoku benchmark – symbolic, visual, and many-solution –, our approach requires a fraction of data and training time of other hybrid methods. On a visual Min-Cut/Max-cut task, it optimizes the regret as well as a Decision-Focused-Learning regret-dedicated loss. Finally, it efficiently learns the energy optimisation formulation of the large real-world problem of designing proteins.
@@ 1. Introduction
In recent years, several hybrid neural architectures have been proposed that integrate discrete reasoning or optimisation within neural network architectures. The motivations for these works are varied. For discrete optimisation and reasoning experts, the main motivation is to deal with uncertainty proactively. Available data is leveraged to produce a solution (decision), whose quality approaches that achievable with a full knowledge of the problem to solve. This motivation has led to the introduction of Decision-Focused Learning (DFL) (mandi2024DFL, 44). Such hybrid architectures are also of interest for users who have a dataset of discrete structured objects of interest (sequences, trees, graphs,…) with associated correlated features and want to be able to predict or generate similar structured objects given only their features. Finally, architectures combining discrete optimisation and learning fit within the neuro-symbolic framework. This framework gathers approaches capable of processing natural inputs – such as images, text or molecules – while simultaneously exhibiting extensive logical reasoning capabilities. Given the accumulating evidence that large language models are limited in their ability to perform logical reasoning (DBLP:conf/icml/BachmannN24, 6, 75, 39, 76), neuro-symbolic systems are a promising alternative research direction in the quest for sparks-of-AGI (artificial general intelligence) (ye2025beyond, 75).
@@
In machine learning, the problem of predicting complex discrete objects (sequences, trees, graphs, …) has been addressed in the field of Structured Output Prediction (SOP): given a training set of pairs $(\omega^{\ell},\mathbf{y}^{\ell})$, where $\omega^{\ell}$ is a suitable input data and $\mathbf{y}^{\ell}$ is a corresponding structured output, the aim is to predict a new high-quality structured output $\mathbf{y}$ from new input data $\omega$. In this setting, $\mathbf{y}$ can be predicted by arbitrary methods. In this paper, we are more specifically interested in architectures where $\mathbf{y}$ is an optimal solution of a predicted combinatorial optimisation (CO) problem instance with unknown objective function $F$ with parameters $\mathbf{c}$ and possibly unknown constraints $\mathcal{C}$. In a generative context, without retraining, the predicted CO instance can be arbitrarily constrained to produce optimal $\mathbf{y}$ that satisfy additional properties. Alternatively, additional objectives can be considered to produce optimal compromises.
@@
This setting raises several challenges. First, the discrete nature of variables defines zero gradients almost everywhere, preventing direct learning by gradient-based optimisation. Second, predicting optimal solutions of CO problems using exact solvers or heuristic algorithms may be computationally demanding and automatic differentiation (autograd) from DL libraries is not applicable. The computational complexity is specifically problematic at training time, with massive training sets that may contain large NP-hard problem instances. We designed our approach to scale to such demanding situations for broad applicability. Our approach can also deal with partially observable settings, where observed solutions $\mathbf{y}^{\ell}$ may contain the values of only a subset of all variables.
@@
Our contributions are threefold: we provide an extensive review of the literature, we introduce a neuro-symbolic architecture combining arbitrary deep learning layers with a final discrete NP-hard Graphical Model (GM) reasoning layer, and propose a new loss function that efficiently deals with logical information presented as constraints. We use discrete GMs (cooper2020graphical, 16) as the reasoning language because GMs have been used to represent both Boolean functions or constraints (in propositional logic, constraint networks) and numerical functions or objectives (in partial weighted Max-SAT, cost function networks). For learning, a probabilistic interpretation of such models is available in stochastic GMs such as Markov Random Fields (MRFs), where infinite costs represent zero probabilities (infeasibility).
@@
We exploit this probabilistic interpretation and target the optimisation of the negative loglikelihood (NLL) of the training set. Because of its intractability, we use an asymptotically consistent and scalable approximation of the NLL known as the Negative Pseudo-LogLikelihood (NPLL) (besag1975statistical, 9). We observe that the NPLL may become incapable of estimating many large costs (difficultGM, 49) and, therefore, of learning constraints. We analyze the reasons for this limitation and propose the E-PLL (Emmental Pseudo-LogLikelihood), which combines NPLL scalability with the ability to learn constraints. At inference, the neural architecture receives natural inputs and outputs a fully-parameterized discrete GM. This model can then be solved using any GM optimisation solver to produce a solution.
@@
With this differentiable informative loss, our architecture efficiently learns constraints from natural inputs to, e.g., solve the Visual Sudoku (satnet, 70, 13). Thanks to its ability to deal with unobserved variables, it can solve the associated symbol grounding problem without cheating (assessingSATNet, 15). The use of an exact prover during inference provides 100% accuracy on hard symbolic Sudoku instances, even with small training sets. The same accuracy is obtained on a many-solution Sudoku data set (1ofMany, 51) and more complex Futoshiki grids (csen2024list, 62). To show that our approach applies in a DFL context where constraints are known, we also tackle Min-Cut and Max-Cut instances where edge weights are represented as images. Finally, as a certificate of scalability, we train our architecture on a set of Protein Design instances with more than 1,000 variables with domain-size $20$.
@@ 2. Related Works
Inverse optimisation and constraint acquisition.
The idea of learning the parameters of optimisation problems from a set of solutions has been explored in various fields. Classical inverse optimisation (chan2023inverse, 14) starts from a dataset of solutions $(\mathbf{y}^{\ell})$, each being an exact optimum $\mathbf{y}^{\ell}\in\operatorname*{arg\,min}F(\mathbf{y},\mathbf{c})$ with unknown parameters $\mathbf{c}$, that must be identified. Similarly, in constraint programming, constraint acquisition starts from sets of solutions, and possibly non-solutions, of an unknown constraint satisfaction (beldiceanu2016modelseeker, 7, 10) or satisfiability (DeRaedt_context, 35) problem that needs to be identified. Compared to our setting, these approaches start from exact solutions/non solutions and cannot exploit arbitrary observable features $\omega$ such as Sudoku images. In data-driven inverse optimisation (chan2023inverse, 14), the optimisation problem parameters $\mathbf{c}$ are split into observable and unobservable parameters $\mathbf{c}=\mathbf{c}_{o}\cup\mathbf{c}_{u}$. Starting from a dataset of pairs $(\mathbf{y}^{\ell},\mathbf{c}_{o}^{\ell})$, the aim is to predict $\mathbf{c}_{u}$ so that on average, empirically, each $\mathbf{y}^{\ell}$ is close, in distance or cost, to an optimal solution of $F(\mathbf{y},\mathbf{c}_{o}^{\ell}\cup\mathbf{c}_{o})$. This can be seen as a specific case of our setting, where the observed data $\omega$ is the exact observation of a fraction of the parameters instead of features known to be correlated with the complete problem definition (i.e., both constraints and objective). However, existing approaches are dedicated to continuous (linear, convex, or conic) optimisation problems.
In its ability to learn constraints, our work is more closely related to the work of (li2023neuro, 36) that combines neural networks with the MaxSAT Z3 solver (de2008z3, 19): from the same input, they predict discrete constraints, by solving a continuous optimisation problem at training time. The first difference with our work lies in the target constraint language used. While we use arbitrary pairwise finite domain constraints, they use “unit” (with 0/1 coefficients) cardinality constraints on Boolean variables. These constraints can have arbitrary arity but take a very restricted form, which may require the introduction of additional latent variables (li2023neuro, 36, § 6). The number of such constraints growing exponentially with the number of variables, a maximum number of learnable constraints, as well as the number of latent variables used, must be set a priori, adding to many hyper-parameters. A second difference lies in the theoretical scalability of both methods: we use an efficient convex loss while they rely on a “Difference of Convex” (DC) programming formulation. While defined on continuous variables, DC is NP-hard in general. Beyond the learning of logical constraints, our loss is also able to learn an objective function, which is required in the context of Decision-focused learning or Protein Design.
@@
Deep learning has been combined with logic programming into several tools such as DeepProbLog (manhaeve2018deepproblog, 45) or NeurASP (NeurASP, 74), where deep learning is in charge of perception and Prolog/ASP is in charge of reasoning. These systems are mostly equipped with the ability of back-propagating the semantic loss (pmlr-v80-xu18h, 72), allowing for training neural nets using logical information, a task that differs from the task considered here.
@@ 2.1. Decision-Focused Learning
In many discrete optimisation under uncertainty problems, the target output $\mathbf{y}$ is the solution of a CO problem defined by a fixed set of constraints $\mathcal{C}$ and an objective function $F$ with unknown parameters $\mathbf{c}$. The parameters $\mathbf{c}$ can be predicted from historic data of features $\omega$, using a training set of pairs $(\omega^{\ell},\mathbf{c}^{\ell})$. In this setting, Predict-then-Optimize – or Prediction-focused-learning – methods (mandi2024DFL, 44) learn to directly predict parameters $\hat{\mathbf{c}}$ from input features $\omega$ with a typical regression loss such as MSE, regardless of the downstream optimisation task. At inference, the fixed discrete model is solved with the predicted $\hat{\mathbf{c}}$. Instead, DFL approaches aim to predict parameters leading to “good” decisions that empirically optimize the average regret, the loss in objective caused by optimizing with $\hat{\mathbf{c}}$ instead of the true parameters $\mathbf{c}$.
DFL differs from our setting by the very nature of the training data. Still, since $\mathcal{C}$ is fixed, DFL training-set pairs $(\omega^{\ell},\mathbf{c}^{\ell})$ can be transformed into pairs $(\omega^{\ell},\mathbf{y}^{\ell})$ where $\mathbf{y}^{\ell}$ satisfies the constraints in $\mathcal{C}$ and optimizes $\mathbf{c}^{\ell}$, just by solving the optimisation problem defined by $\mathcal{C}$ and $\mathbf{c}^{\ell}$.
Thus, methods designed for SOP directly apply to DFL settings and have been de facto considered as DFL tools in recent reviews (mandi2024DFL, 44). Existing methods have been classified in four families (mandi2024DFL, 44): analytical differentiation of optimisation mappings, analytical smoothing of optimisation mappings, smoothing by random perturbations, and differentiation of surrogate loss functions. Differentiable optimisation layers have mostly been developed for convex or conic continuous optimisation problems. They can be applied to discrete optimisation problems after proper smoothing. This is, for instance, the case of SAT-Net (satnet, 70), which relaxes the discrete MaxSAT problem into a convex semi-definite programming problem. SAT-Net has been used to learn how to solve easy Sudoku and Visual Sudoku (from grid images). It has been pointed out that SAT-Net was ineffective if labels for observed digit images were not provided in the supervised training set (assessingSATNet, 15).
@@
Analytical smoothing is also needed for learning linear programming (LPs) problems because of the intrinsically discrete nature of the polytope’s vertices. One standard smoothing is the addition of an $L_{2}$ regularization on solutions (Wilder_Dilkina_Tambe_2019, 71, 52). For integer LPs (ILPs), it is possible to relax them to LPs (Wilder_Dilkina_Tambe_2019, 71), possibly with cutting plane strengthening (MIPaaL2020, 24) for a better polytope approximation. Discrete (Boolean/One Hot Encoded) pairwise GMs, the formalism we use in this paper, are intrinsically quadratic in their discrete decision variables and linear in their parameters. They can be optimized by reduction to an 01LP, requiring the introduction of a quadratic number of variables. This 01LP model, relaxed to continuous variables, is known as the GM’s “local polytope” (cooper2020graphical, 16). Existing approaches for learning ILP could, therefore, be used to learn GMs. Due to their size, the associated (I)LP reductions are challenging to solve (hurley2016multi, 29, 4).
@@
Besides smoothing by relaxation or regularization, e.g.  in Fenchel-Young losses (blondel2020learning, 12), smoothing can be achieved by random perturbations (Berthet, 8). These approaches rely on a probabilistic interpretation of the linear objective function using maximum-entropy exponential distributions. For discrete models, assuming minimisation, this usually means Gibbs/Boltzmann distributions with the probability $p(\mathbf{y})\propto\exp(-\langle\mathbf{c},\mathbf{y}\rangle)$. In arbitrary discrete settings, the resulting probability distributions suffer from the intractability of computing the normalizing constant $Z$. This can sometimes be simplified using the Gumbel trick (huijben2022review, 28), which allows for estimating $Z$ using a series of calls to an optimisation oracle on a Gumbel-perturbed objective parameters $\mathbf{c}$ (iMLE, 53) or using other perturbations (Berthet, 8). These approaches can be interpreted as Fenchel-Young losses (blondel2020learning, 12, 8), where the perturbed solutions provide the regularization. The stochastic variants of GMs (koller2009probabilistic, 33, 16) we rely on, Markov Random Fields, are decomposable log-linear models that rely on the same family of exponential distributions. Their log-likelihood is intractable for the same reason. It is known to be a Fenchel-Young loss (blondel2020learning, 12), often referred to as the CRF (Conditional Random Field) loss. We instead use and improve a tractable approximation of log-likelihood that exploits the decomposability of GMs.
@@
These approaches try to optimize regret indirectly. The SPO+ loss (SPO, 23) is an upper bound of regret with a gradient that resembles perturbed optimizers’ gradients. A recent related approach, based on Noise Contrastive Estimation (Mulamba21, 50), exploits an incrementally built pool of feasible solutions in a probabilistic interpretation, as above.